OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (mimi_x3):

\[y = mx + b \] Sub x = 2 , y = -4 , and m = 3/5 into the equation You will get: \[-4 = \frac{3}{5} (2) + b \] Then solve for b, for the equation of the line.

OpenStudy (mimi_x3):

Is this another question ? :/